OpenStudy (anonymous):

simplify this expression cos(theta, the 0with a line through it)/tan(theta)cot(theta) A)sin(theta) B)cos(theta) C)csc(theta) D)sec(theta)

OpenStudy (anonymous):

lol 0 with a line \(\theta\) right?

OpenStudy (anonymous):

the variable is unimportant

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

rewrite all in terms of sine and cosine

OpenStudy (anonymous):

i didn't know how to put it in but that was the symbol

OpenStudy (anonymous):

no problem lets dispense with the trig all together

OpenStudy (anonymous):

call \(\sin(\theta)=b, \cos(\theta)=a\) then this is the same as \[a\times \frac{b}{a}\times \frac{a}{b}\]

OpenStudy (anonymous):

or even easier, know that cotangent is the reciprocal of tangent so \[\tan(x)\cot(x)=1\]

OpenStudy (anonymous):

i'm a little lost

OpenStudy (anonymous):

lets go slow then

OpenStudy (anonymous):

so it could be cos*sin/cos*cos/sin?

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

and when you are done cancelling you are left only with \(\cos(\theta)\)

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw